OpenStudy (anonymous):

If a is constant, evaluate the integral below. Give your answer using words or symbols. Remember that ^ can be used for exponents. x^4 dx x is from 1 to a help??

OpenStudy (turingtest):

simply apply the basic rule\[\int x^ndx=\frac{x^{n+1}}{n+1}\]and evaluate

OpenStudy (anonymous):

so what would i plug in for n?

OpenStudy (anonymous):

\[ \int_1^a x^4 dx=\left.\frac{x^5}{5}\right|_1^a\]

OpenStudy (anonymous):

oh...that's it?

OpenStudy (anonymous):

You have to plug in the 'a' and '1' and subtract.